a child of mass 40 kg slides down a water slide that is 15 m high. at the bottom her velocity is 9.7 m/s. approximately how much energy was "lost" due to friction?


1
Expert's answer
2022-01-30T13:51:57-0500

Solution

Mass of child m=40kg

Height h=15m

Velocity at bottom

v=9.7m/s

So energy at height h

U=mgh

="40\times9.81\times15"

"=5886J"

Energy at bottom

E="0.5mv^2"

"=0.5\times40\times(9.7) ^2\\=1881.8J"

Loss of energy due to friction

"=U-E\\=5886-1881.8\\=4004.2J"